A stone is thrown vertically upwards from the ground with an initial velocity of 20ms-1.

a)What is the maximum height reached?

b)How long does the ball take to reach this maximum height?

c)How long does it take for the ball to fall to the ground again from the maximum height?


1
Expert's answer
2022-06-01T12:51:47-0400

(a)

"h_{\max}=\frac{v_0^2}{2g}=\frac{20^2}{2*9.8}=20\:\rm m"

(b)

"t=\frac{v_0}{g}=\frac{20}{9.8}=2.0\:\rm s"

(c)

"t=\sqrt{\frac{2h_{\max}}{g}}=\sqrt{\frac{2*20}{9.8}}=2.0\:\rm s"